Circular No. 9049

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V5582 SAGITTARII G. Sun, Wenzhou, Zhejiang, China; and X. Gao, Urumqi, Xinjiang,

China, report the discovery of a possible nova (mag ≈ 11.5) on several 60-s survey images (limiting mag ∼ 13.5) taken by Gao in the course of their nova survey at Mt. Nanshan on Feb. 23.947–23.963 UT using a Canon EOS 350D camera (+ 135-mm f/2 lens). Their images taken on Apr. 28 by Sun with a 1-m f/8 reflector at Weihai yield the following position for the variable: α = 17^{h}45^{m}05^{s}.40, δ = −20^{o}03′21′′.5 (equinox 2000.0). Nothing was visible at this location on images taken by the discoverers on 2008 Aug. 23, Sept. 4, 25, and Oct. 3 (limiting mag presumably also ∼ 13.5); they report that nothing is visible at this position on Digitized Sky Survey images from 1950 June 19 (limiting red mag 20.0), 1980 Aug. 2 (limiting infrared mag 19.5), and 1991 Aug. 2 (limiting red mag 20.8). Additional available magnitudes from Sun and Gao for the variable: Feb. 27.960, 12.1; 28.968, 12.0; Mar. 2.963, 13.0; 3.964, 12.8; 6.968, 13.1; 22.949, 11.6; 26.950, 12.1; 27.947, 12.5; Apr. 28.837, 13.2.

Following posting on the Central Bureau’s unconfirmed-objects webpage, other observers have reported their observations of this variable. K. Nishiyama (Kurume, Fukuoka, Japan) and F. Kabashima (Miyaki, Saga, Japan) report (via S. Nakano, K. Kinugasa, and H. Yamaoka) the following magnitudes from their CCD frames: Feb. 20.855, 10.9; 25.847, 10.4; 28.859, 11.4; Mar. 14.860, 13.2; May 19.665, 14.0; 22.728, 13.7. From a CCD image taken with a 40-cm reflector, Nishiyama and Kabashima measured position end figures 05^{s}.40, 21′′.7. G. Sostero, E. Guido, and P. Camilleri write that they obtained position end figures 05^{s}.42, 22′′.0 and magnitudes R = 12.2, B = 13.8 from images obtained remotely on Mar. 13.44 with a 25-cm f/3.4 reflector at the GRAS Observatory near Mayhill, NM, U.S.A.; comparison with an Anglo-Australian Observatory Schmidt red plate (limiting magnitude ∼ 20), obtained on 1991 Aug. 2, shows an extremely faint object at position end figures of 05^{s}.38, 23′′.4 at the threshold of the plate. Additional details are given on CBET 1816.

K. Kinugasa, S. Honda, and O. Hashimoto, Gunma Astronomical Observatory (GAO); and Y. Takeda, National Astronomical Observatory of Japan, obtained a low-resolution spectrum (range 400–800 nm, resolution ∼ 500) of this object using the GAO 1.5-m telescope (+ GLOWS) on May 26.7 UT, which shows strong emission lines of Balmer series, [O III], [N II], and He I — suggesting that the object is a classical nova well past maximum. N. Samus adds that the GCVS team has assigned the designation V5582 Sgr to this nova.

2009 May 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green